OpenStudy (anonymous):

Can some one please explain Quadratic Equations give me some examples please going to do a test on it and im lost

OpenStudy (anonymous):

A quadratic equation is a polynomial that has a degree of 2. For example... \[x^2 + 7x + 6 = 0\] You can use the quadratic formula to solve for the roots of it, or look at it and figure out the sum of the roots and their product. In this case, their sum is 7 and their product is 6. \[(x + 1)(x + 6) = 0\] 6 + 1 = 7 6 * 1 = 6

OpenStudy (lilg132):

x = -1 and x = -6

OpenStudy (lilg132):

or use the quadratic formula \[x = {-b \pm \sqrt{b ^{2} - 4ac} \over 2a}\]

OpenStudy (lilg132):

remember to click good answer